So Im doing solving equations with the variable on both sides of =. I can do that easily but when fractions on both sides are involved it gets a little trickier but I can still do it by multiplying each side by the lcd. The question I'm asking is when there are fractions on both side of the equations with the variables in them, for example. x/3 + 3 = 2/6x + 6 . This isnt a real equation (or it might be) but its just for example. So how do I find solve the equations with the variables on both sides of = with fractions? I don't know if I can ask more wordy questions on this sight but im confused and want to do math better.
Found 2 solutions by Boreal, MathTherapy:
Answer by Boreal(15235)

Multiply everything by the LCD
So if you have the equation above
x/3 + 3 = 2/6x + 6 ,
first, make life easier and move the 3 over
(x/3)=(2/6x)+3
Now multiply everything by 6x
2x(x)=2+18x
2x^2-18x+2=0, divide by 2
x^2-9x+1=0, and here you would solve using the quadratic formula.

(x/2)+(6/2x)=3
multiply everything by 2x, the LCD.
x^2+6=6x, move things over and solve.

Answer by MathTherapy(10551)

So Im doing solving equations with the variable on both sides of =. I can do that easily but when fractions on both sides are involved it gets a little trickier but I can still do it by multiplying each side by the lcd. The question I'm asking is when there are fractions on both side of the equations with the variables in them, for example. x/3 + 3 = 2/6x + 6 . This isnt a real equation (or it might be) but its just for example. So how do I find solve the equations with the variables on both sides of = with fractions? I don't know if I can ask more wordy questions on this sight but im confused and want to do math better.
First of all, you need to use parentheses!
Based on what you presented, it can be either of the following:
If it's the former, you'd need to multiply the equation by the LCD, 6 to get: 2x + 18 = 2x + 36, which by the way, has NO solutions.
However, it's the latter, you'd need to multiply the equation by the LCD, 6x to get: 2x2 + 18x = 2 + 36x.